# Copositive criteria for a two-component dark matter model

### v1

@@ Abstract
We consider a two-component scalar dark matter model in this work, where the scalars are stabilized by extra $Z_{2}\times Z^{\prime}_{2}$ symmetry. To guarantee the stability of the vacuum, we consider the copositive criteria and different choices of the signs of the couplings will contribute to different viable parameter spaces. Based on the copositive criteria, we systematically analyze the possible conditions and pick up 17 different cases with fixing some parameters to be negative. We randomly scan the parameter space under dark matter relic density constraint and direct detection constraint and focus on the quartic couplings with $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$. The shapes of the viable parameter space for $|\lambda_{14}|-|\lambda_{24}|$ are almost similar among these cases, while for $|\lambda_{13}|$ and $|\lambda_{23}|$, the larger value are excluded as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ due to the large interference effect of the $2\to 2$ processes.
1 Introduction
@@
The Standard Model (SM) of particle physics has been remarkably successful in describing the fundamental particles and their interactions. However, the astrophysical observations indicate the existence of dark matter, which constitutes approximately $26.5\%$ of the universe’s mass-energy content Planck:2015fie . The existence of dark matter can be inferred from its gravitational effects on the universe’s large-scale structure, yet its particle nature remains elusive. From the view of particle physics, dark matter can be assumed to be some stable particle carrying no SM gauge group quantum number.
@@
One of the most popular and attractive dark matter candidates is weakly interacting massive particles(WIMPs)Qi:2021rhh ; Qi:2022fzs ; Qi:2021rpa , where the observed dark matter relic density is generated via the so-called “Freeze-out” mechanismRoszkowski:2017nbc . Among the different WIMP scenarios, the singlet scalar dark matter model is the simplest where a new singlet scalar as dark matter is introduced to the SM. The singlet scalar is stabilized by extra $Z_{2}$ symmetry, and there are only two free model parameters in the model with dark matter mass and the quartic coupling of dark matter with SM Higgs. According to Higgs invisible decay search at LHCATLAS:2016neq , dark matter relic density, direct detection constraint He:2016mls ; Escudero:2016gzx and indirect detection constraint arising from gamma-ray spectrum induced by DM annihilation at Fermi-LAT Fermi-LAT:2015kyq ; Fermi-LAT:2015att and HESSHESS:2013rld , the viable dark matter mass region is limited stringently with resonant mass being about half of the SM Higgs mass and high mass region larger than about 3 TeV Wu:2016mbe . The current direct detection experiments have not yielded a clear signal, imposing stringent constraints on DM-nucleon interactions, so the WIMP scenarios are facing a serious crisis.
@@
One possible solution to the problem is the multi-component dark matter models, which include two or more dark matter species. In the multi-component dark matter models, it is possible that one of the dark matter candidates has a large elastic cross-section on nuclei but forms only a subdominant DM component leading to a suppressed signal Belanger:2022qxt therefore escapes from the direct detection constraint. There is no preference to assume that there is only one dark matter candidate in the universe, and discussion about multi-component dark matter can be found in Coleppa:2023vfh ; Zurek:2008qg ; Costa:2022lpy and so on. Particularly, for the multi-component scalar dark matter model, one can introduce extra symmetry such as $Z_{2}\times Z_{2}^{\prime}$hattacharya:2019fgs , $Z_{5}$ symmetry Belanger:2020hyh ; Qi:2024uiz ; Qi:2023egb ,$Z_{7}$ symmetry Belanger:2022esk or other $Z_{N}$ symmetry Bhattacharya:2024nla ; Yaguna:2019cvp ; Yaguna:2021vhb to ensure the stability of dark matter particles.
@@
Discussion about copositive criteria and dark matter can be found in DuttaBanik:2020jrj ; Kannike:2012pe ; Chakrabortty:2013mha ; Arhrib:2011uy , and in this work, we consider a simple two-component scalar dark matter model and discuss the copositive criteria on the model. We introduce two singlet scalars $S_{1}$ and $S_{2}$ as dark matter to the SM, which are stabilized by $Z_{2}\times Z_{2}^{\prime}$ symmetry. In addition, another singlet scalar $S_{3}$ with non-zero vacuum expectation value is also introduced to the model, therefore, a $4\times 4$ copositive matrix of the quartic couplings can be constructed in the model. Note that the copositive criteria of the rank $\leqslant 3$ matrix can be given concretely, while for the higher rank, related discussions about the models are always ignored for their complexity. In this paper, we analyze the copositive criteria of the model systematically. Generally speaking, there are eight cases of the copositive criteria for the rank 4 matrix, and one can have more complex cases when considering the signs of the couplings of the model. We pick up 17 different cases from the point of exchange symmetry for simplicity and estimate the viable parameter space within the dark matter relic density constraint and direct detection constraint. The signs of the couplings will not only determine the copositive criteria but also make a difference in the cross section of $2\to 2$ processes related to the dark matter annihilation into other scalars due to the interference effect. Within the copositive criteria, we randomly scan the chosen parameter space and focus on the quartic couplings with $\lambda_{13}$,$\lambda_{23}$,$\lambda_{14}$ and $\lambda_{24}$. We found the shapes of the viable parameter space for $|\lambda_{14}|-|\lambda_{24}|$ are almost similar among these cases, while for $|\lambda_{13}|$ and $|\lambda_{23}|$, the larger value are excluded as long as $\lambda_{13}\leqslant 0$ and $\lambda_{23}\leqslant 0$ due to the large interference effect.
@@
The paper is organized as follows. We give the two-component scalar dark matter in Sec.2. In Sec.3, we discuss the copositivity of the symmetric matrix with eight different cases. In Sec.4, we consider the dark matter phenomenology of the model including relic density and direct detection. In Sec.5, we estimate the parameter space from the view of copositive criteria, we pick up 17 different permutations of the indexes and give the viable parameter space satisfying relic density and direct detection constraints. Finally, we give a summary in Sec.6.
2 Model description
@@
In this section, we consider a simple two-component scalar dark matter model with $Z_{2}\times Z_{2}^{\prime}$ symmetry by introducing three singlet scalars $S_{1,2,3}$, where the corresponding charges for dark matter particles $S_{1}$ and $S_{2}$ are $(-1,1)$ and $(1,-1)$ and $S_{3}$ has non-zero vacuum expectation value (vev), while the SM particles carry $(1,1)$ charge. In addition, we consider another $Z_{2}^{\prime\prime}$ symmetry for $S_{3}$ for simplicity, which can arise from a spontaneously broken $U(1)$ gauge symmetry. Therefore, the potential part of the scalars can be given by:
$\mathcal{V}$ $=\frac{1}{2}M_{1}^{2}S_{1}^{2}+\frac{1}{2}M_{2}^{2}S_{2}^{2}-\mu_{3}^{2}|S_{3}|^{2}-\mu_{H}^{2}|H|^{2}+\lambda_{H}|H|^{4}+\frac{1}{4}\lambda_{11}S_{1}^{4}+\frac{1}{4}\lambda_{22}S_{2}^{4}+$
$\lambda_{33}|S_{3}|^{4}+\lambda_{12}S_{1}^{2}S_{2}^{2}+\lambda_{13}S_{1}^{2}|S_{3}|^{2}+\lambda_{14}S_{1}^{2}|H|^{2}+\lambda_{23}S_{2}^{2}|S_{3}|^{2}+\lambda_{24}S_{2}^{2}|H|^{2}$
$+\lambda_{34}|S_{3}|^{2}|H|^{2}$ (1)
@@
where $H$ is the SM Higgs doublet. Under unitarity gauge, $H$ and $S_{3}$ can be expressed with:
$H=\left(\begin{array}[]{c}0\\ \frac{v_{0}+h}{\sqrt{2}}\end{array}\right)\,,\quad S_{3}=s_{3}+v_{1}\,,\quad$ (2)
@@
$m_{1}^{2}=M_{1}^{2}+2\lambda_{13}v_{1}^{2}+\lambda_{14}v_{0}^{2},m_{2}^{2}=M_{2}^{2}+2\lambda_{23}v_{1}^{2}+\lambda_{24}v_{0}^{2}$ (3)
@@
$m^{2}_{h_{2}}$ $=\lambda_{H}v_{0}^{2}+\lambda_{33}v_{1}^{2}+\sqrt{(\lambda_{H}v_{0}^{2}-\lambda_{33}v_{1}^{2})^{2}+(\lambda_{34}v_{0}v_{1})^{2}}$ (7)
@@
$\tan 2\theta=\frac{\lambda_{34}v_{0}v_{1}}{\lambda_{33}v_{1}^{2}-\lambda_{H}v_{0}^{2}}$ (9)
@@
$\lambda_{34}$ $=\frac{\sin 2\theta(m_{h_{2}}^{2}-m_{h_{1}}^{2})}{2v_{0}v_{1}}$
3 Copositivity of symmetric matrix
@@
In the following discussion, we use $\lambda_{44}$ to represent $\lambda_{H}$ to make our discussion more coherent. For this symmetric matrix of order four, we should convict eight different cases depending on the sign distributions of the off-diagonal elements to determine whether this matrix is copositive. In these cases, all the diagonal elements should be positive as a generic condition, and the eight cases are given as follows Chakrabortty:2013mha :
@@
4 Dark matter phenomenology
In our model, there are two scalar dark matter particles stabilized by $Z_{2}\times Z_{2}^{\prime}$, which are related to SM particles with Higgs-portal interactions. The current observed dark matter relic density given by the Planck collaboration is $\Omega_{DM}h^{2}=0.1198\pm 0.0012$ Belanger:2020gnr , and we consider dark matter production in our model to be generated with the “Freeze-out” mechanism. Both $S_{1}$ and $S_{2}$ will contribute to dark matter relic density and the Boltzmann equations for the number density of $S_{1}$ and $S_{2}$ are given as follows:
$\frac{dn_{1}}{dt}+3Hn_{1}$ $=$ $-\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to XX}(n_{1}^{2}-{\bar{n_{1}}}^{2})-\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to h_{1,2}h_{1,2}}(n_{1}^{2}-{\bar{n_{1}}}^{2})$
-- $\langle\sigma v\rangle^{S_{1}S_{1}^{*}\to S_{2}S_{2}^{*}}(n_{1}^{2}-n_{2}^{2}\frac{\bar{n_{1}}^{2}}{\bar{n_{2}}^{2}})$
$\frac{dn_{2}}{dt}+3Hn_{2}$ $=$ $-\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to XX}(n_{2}^{2}-{\bar{n_{2}}}^{2})-\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to h_{1,2}h_{1,2}}(n_{2}^{2}-{\bar{n_{2}}}^{2})$ (16)
-- $\langle\sigma v\rangle^{S_{2}S_{2}^{*}\to S_{1}S_{1}^{*}}(n_{2}^{2}-n_{1}^{2}\frac{\bar{n_{2}}^{2}}{\bar{n_{1}}^{2}})$
where $n_{1}(n_{2})$ is the number density of $S_{1}(S_{2})$, $\bar{n_{1}}(\bar{n_{2}})$ is the number density in thermal equilibrium, $H$ is the Hubble expansion rate of the Universe, $X$ denotes SM particles and $\langle\sigma v\rangle$ is the thermally averaged annihilation cross section. To calculate the DM relic density numerically we use the micrOMGEAs 5.0.6 package Alguero:2022inz , in which the model has been implemented through the FeynRules package Alloul:2013bka .
On the other hand, dark matter can scatter off the nuclei with the Higgs-mediated t-channel processes, which should be constrained by direct detection results. The effective Lagrangian related to DM-quark elastic scattering can be given by Basak:2021tnj :
$\mathcal{L}_{q,eff}=\sum_{S=S_{1},S_{2}}-\frac{m_{q}}{2v_{0}}(\frac{\mathcal{C}_{h_{1}SS}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}SS}}{m_{h_{2}}^{2}})SS\bar{q}q$ (17)

### v4

@@ Abstract
In this work, we investigate a two-component scalar dark matter framework featuring two singlet scalar fields as dark matter candidates. To ensure vacuum stability, we employ copositive criteria for the scalar potential. Guided by direct detection constraints, we analyze four distinct copositive scenarios characterized by specific negative parameter configurations. A comprehensive parameter space scan is performed under joint constraints from observed dark matter relic density and direct detection experiments. Different signs of the couplings not only correspond to different copositive criteria but can also contribute to different parameter spaces due to the interference effect. For the 4 different cases, the allowed values of quartic couplings are different but all demand that new Higgs plays a dominant role in determining dark matter relic density within the viable parameter space.
I Introduction
@@
The Standard Model (SM) of particle physics has been remarkably successful in describing the fundamental particles and their interactions. However, the astrophysical observations indicate the existence of dark matter, which constitutes approximately $26.5\%$ of the universe’s mass-energy content Ade and others (2016). The existence of dark matter can be inferred from its gravitational effects on the universe’s large-scale structure, yet its particle nature remains elusive. From the view of particle physics, dark matter can be assumed to be some stable particle carrying no SM gauge group quantum number.
@@
One of the most popular and attractive dark matter candidates is weakly interacting massive particles(WIMPs)Qi et al. (2022); Qi and Sun (2023b, c), where the observed dark matter relic density is generated via the so-called “Freeze-out” mechanismRoszkowski et al. (2018). Among the different WIMP scenarios, the singlet scalar dark matter model is the simplest where a new singlet scalar as dark matter is introduced to the SM. The singlet scalar is stabilized by extra $Z_{2}$ symmetry, and there are only two free model parameters in the model with dark matter mass and the quartic coupling of dark matter with SM Higgs. According to Higgs invisible decay search at LHCAad and others (2016), dark matter relic density, direct detection constraint He and Tandean (2016); Escudero et al. (2016) and indirect detection constraint arising from gamma-ray spectrum induced by DM annihilation at Fermi-LAT Ackermann and others (2015b, a) and HESSAbramowski and others (2013), the viable dark matter mass region is limited stringently with resonant mass being about half of the SM Higgs mass and high mass region larger than about 3 TeV Wu and Zheng (2017). The current direct detection experiments have not yielded a clear signal, imposing stringent constraints on DM-nucleon interactions, so the WIMP scenarios are facing a serious crisis.
@@
One possible solution to the problem is the multi-component dark matter models, which include two or more dark matter species. In the multi-component dark matter models, it is possible that one of the dark matter candidates has a large elastic cross-section on nuclei but forms only a subdominant DM component leading to a suppressed signal Belanger et al. (2022) therefore escapes from the direct detection constraint. There is no preference to assume that there is only one dark matter candidate in the universe, and discussion about multi-component dark matter can be found in Coleppa et al. (2024); Zurek (2009); Costa et al. (2022) and so on. Particularly, for the multi-component scalar dark matter model, one can introduce extra symmetry such as $Z_{2}\times Z_{2}^{\prime}$ Bhattacharya et al. (2020) , $Z_{5}$ symmetry Bélanger et al. (2020); Qi and Sun (2024, 2023a) $Z_{7}$ symmetry Bélanger et al. (2023) or other $Z_{N}$ symmetry Bhattacharya et al. (2024); Yaguna and Zapata (2020, 2021) to ensure the stability of dark matter particles.
@@
Discussion about copositive criteria and dark matter can be found in Dutta Banik et al. (2021); Kannike (2012); Chakrabortty et al. (2014); Arhrib et al. (2011); Ghorbani (2021); Habibolahi et al. (2022), and in this work, we consider a simple two-component scalar dark matter model and discuss the copositive criteria on the model. We introduce three singlet scalars $S_{1}$, $S_{2}$, and $S_{3}$ to the SM, where $S_{1}$ as well as $S_{2}$ as dark matter candidates and the singlet scalar $S_{3}$ has non-zero vacuum expectation value. Note that we come to the two component scalar dark matter model Díaz Sáez et al. (2021) when $S_{3}$ is absent, which demands a large $\lambda_{12}$ to evade direct detection and relic abundance constraints outside the Higgs resonance region. With the introduction of $S_{3}$, we have new annihilation channels and the couplings of DM with SM Higgs as well as $\lambda_{12}$ can be smaller to escape from the direct detection constraints with the observed DM relic density constraint. In other word, the existence of $S_{3}$ will contribute to the new dark matter annihilation processes so that we can have a wider parameter space compared with the dark matter models that only involve two singlet scalars. On the other hand, a $4\times 4$ copositive matrix of the quartic couplings can be constructed in the model. Note that the copositive criteria of the rank $\leqslant 3$ matrix can be given concretely, while for the higher rank, related discussions about the models are always ignored for their complexity. In this paper, we analyze the copositive criteria of the model systematically. Generally speaking, there are eight cases of the copositive criteria for the rank 4 matrix, and one can have more complex cases when considering the signs of the couplings of the model. Based on the direct detection constraint, we pick up 4 different cases from the point of exchange symmetry for simplicity and estimate the viable parameter space within the dark matter relic density constraint and direct detection constraint. The signs of the couplings will not only determine the copositive criteria but also make a difference in the cross-section of $2\to 2$ processes related to the dark matter annihilation into other scalars due to the interference effect. We randomly scan the chosen parameter space and focus on the quartic couplings with $\lambda_{13}$$\lambda_{23}$$\lambda_{14}$ and $\lambda_{24}$. For the 4 different cases, the allowed values of $(|\lambda_{13}|,|\lambda_{14}|)$ are different but all demand that new Higgs plays a dominant role in determining dark matter relic density. Within the viable parameter space, copositive criteria can put stringent constraints as long as some coupling is negative.
@@
The paper is organized as follows. We give the two-component scalar dark matter model in Sec.II. In Sec.III, we discuss the copositivity of the symmetric matrix with eight different cases. In Sec.IV, we consider the dark matter phenomenology of the model including relic density and direct detection. In Sec.V, we pick up 4 different cases by fixing some of the quartic parameters to be negative and discuss the effect of copositive constraint on the viable parameter space satisfying relic density and direct detection constraints. Finally, we give a summary in Sec.VI.
II Model description
@@
In this section, we consider a simple two-component scalar dark matter model with $Z_{2}\times Z_{2}^{\prime}\times Z_{2}^{\prime\prime}$ symmetry by introducing three singlet scalars $S_{1,2,3}$, where the corresponding charges for dark matter particles $S_{1}$ and $S_{2}$ are (1,1,1)(-1,1,1) and (1,1,1)(1,-1,1) and $S_{3}$ has non-zero vacuum expectation value (vev) carrying (1,1,1)(1,1,-1) charge, while the SM particles carry $(1,1,1)$ charge. Therefore, the potential part of the scalars can be given by:
$\mathcal{V}$ $=\frac{1}{2}M_{1}^{2}S_{1}^{2}+\frac{1}{2}M_{2}^{2}S_{2}^{2}-\frac{1}{2}\mu_{3}^{2}S_{3}^{2}-\mu_{H}^{2}|H|^{2}+\lambda_{H}|H|^{4}+\frac{1}{4}\lambda_{11}S_{1}^{4}+\frac{1}{4}\lambda_{22}S_{2}^{4}+\frac{1}{4}\lambda_{33}S_{3}^{4}$
$+\lambda_{12}S_{1}^{2}S_{2}^{2}+\lambda_{13}S_{1}^{2}S_{3}^{2}+\lambda_{14}S_{1}^{2}|H|^{2}+\lambda_{23}S_{2}^{2}S_{3}^{2}+\lambda_{24}S_{2}^{2}|H|^{2}+\lambda_{34}S_{3}^{2}|H|^{2}.$ (1)
@@
where $H$ is the SM Higgs doublet. Note that the $Z_{2}^{\prime\prime}$ symmetry is introduced to simplify the scalar potential in this work. Under unitarity gauge, $H$ and $S_{3}$ can be expressed with:
$H=\left(\begin{array}[]{c}0\\ \frac{v_{0}+h}{\sqrt{2}}\end{array}\right)\,,\quad S_{3}=s_{3}+v_{1}\,$$\quad$ (2)
@@
$m_{1}^{2}=M_{1}^{2}+2\lambda_{13}v_{1}^{2}+\lambda_{14}v_{0}^{2},m_{2}^{2}=M_{2}^{2}+2\lambda_{23}v_{1}^{2}+\lambda_{24}v_{0}^{2}.$ (3)
@@
$m^{2}_{h_{2}}$ $=\lambda_{H}v_{0}^{2}+\lambda_{33}v_{1}^{2}+\sqrt{(\lambda_{H}v_{0}^{2}-\lambda_{33}v_{1}^{2})^{2}+(\lambda_{34}v_{0}v_{1})^{2}}.$ (7)
@@
$\tan 2\theta=\frac{\lambda_{34}v_{0}v_{1}}{\lambda_{33}v_{1}^{2}-\lambda_{H}v_{0}^{2}}.$ (9)
@@
$\lambda_{34}$ $=\frac{\sin 2\theta(m_{h_{2}}^{2}-m_{h_{1}}^{2})}{2v_{0}v_{1}}.$
III Copositivity of symmetric matrix
@@
In the following discussion, we use $\lambda_{44}$ to represent $\lambda_{H}$ to make our discussion more coherent. For this symmetric matrix of order four, we should convict eight different cases depending on the sign distributions of the off-diagonal elements to determine whether this matrix is copositive. In these cases, all the diagonal elements should be positive as a generic condition, and the eight cases are given as follows Chakrabortty et al. (2014):
@@
IV Dark matter phenomenology
In our model, there are two scalar dark matter particles stabilized by $Z_{2}\times Z_{2}^{\prime}\times Z_{2}^{\prime\prime}$ symmetry, which are related to SM particles with Higgs-portal interactions. The current observed dark matter relic density given by the Planck collaboration is $\Omega_{DM}h^{2}=0.1198\pm 0.0012$ Belanger et al. (2021), and we consider dark matter production in our model to be generated with the “Freeze-out” mechanism. Both $S_{1}$ and $S_{2}$ will contribute to dark matter relic density and the Boltzmann equations for the number density of $S_{1}$ and $S_{2}$ are given as follows:
$\frac{dn_{1}}{dt}+3Hn_{1}$ $=$ $-\langle\sigma v\rangle^{S_{1}S_{1}\to XX}(n_{1}^{2}-{\bar{n_{1}}}^{2})-\langle\sigma v\rangle^{S_{1}S_{1}\to h_{1,2}h_{1,2}}(n_{1}^{2}-{\bar{n_{1}}}^{2})$
- $\langle\sigma v\rangle^{S_{1}S_{1}\to S_{2}S_{2}}(n_{1}^{2}-n_{2}^{2}\frac{\bar{n_{1}}^{2}}{\bar{n_{2}}^{2}}),$
$\frac{dn_{2}}{dt}+3Hn_{2}$ $=$ $-\langle\sigma v\rangle^{S_{2}S_{2}\to XX}(n_{2}^{2}-{\bar{n_{2}}}^{2})-\langle\sigma v\rangle^{S_{2}S_{2}\to h_{1,2}h_{1,2}}(n_{2}^{2}-{\bar{n_{2}}}^{2})$ (16)
- $\langle\sigma v\rangle^{S_{2}S_{2}\to S_{1}S_{1}}(n_{2}^{2}-n_{1}^{2}\frac{\bar{n_{2}}^{2}}{\bar{n_{1}}^{2}}).\ \ \ \$
where $n_{1}(n_{2})$ is the number density of $S_{1}(S_{2})$, $\bar{n_{1}}(\bar{n_{2}})$ is the number density in thermal equilibrium, $H$ is the Hubble expansion rate of the Universe, $X$ denotes SM particles and $\langle\sigma v\rangle$ is the thermally averaged annihilation cross section. In Fig. 1, we show the Feynman diagrams of dark matter annihilation processes related to new particles. Note that both $S_{1}$ and $S_{2}$ can be dominant in DM relic density due to the exchange symmetry, which depends on the mass heirarchy between $S_{1}$ and $S_{2}$, and the lighter one will be dominant since heavy component can annihilate into the lighter component. To calculate the DM relic density numerically we use the micrOMGEAs 5.0.6 package Alguero et al. (2022), in which the model has been implemented through the FeynRules package Alloul et al. (2014).
Figure 1: Feynman diagrams of dark matter annihilation processes related to new particles.
On the other hand, dark matter can scatter off the nuclei with the Higgs-mediated t-channel processes, which should be constrained by direct detection results. The effective Lagrangian related to DM-quark elastic scattering can be given by Basak et al. (2021):
$\mathcal{L}_{q,eff}=\sum_{S=S_{1},S_{2}}-\frac{m_{q}}{2v_{0}}(\frac{\mathcal{C}_{h_{1}SS}}{m_{h_{1}}^{2}}+\frac{\mathcal{C}_{h_{2}SS}}{m_{h_{2}}^{2}})SS\bar{q}q.$ (17)